A particle is projected horizontally from the top of the cliff of height ‘H’ with a velocity of √(2gH). The radius of curvature of the trajectory at the instant of projection will be?

Grade:9

1 Answers

Eshan
5 years ago
Speed of projection=\sqrt{2gH}

The acceleration of the body at this moment=g

Assuming the motion to be circular momentartily, the acceleration of the body would be given as

g=\dfrac{v^2}{R}
\implies R=\dfrac{v^2}{g}=2H